# GEMM-GS: Accelerating 3D Gaussian Splatting on Tensor Cores with GEMM-Compatible Blending

### 02 Apr 2026 v1

@@ 2. Background
where $\mathbf{C}_{j}$ is the rendered color of pixel $p_{j}$, and $T_{i}$ represents the accumulated transmittance up to Gaussian $i$.
@@ 8    Fetch data of Gaussian $\mathcal{G}_{i}$:
10      covariance matrix: $\Sigma_{i}=\begin{bmatrix}A_{i}&B_{i}\\ B_{i}&C_{i}\end{bmatrix}$
@@ 3.1. Calculation Form of $\alpha$’s Exponential
where $\mathbf{x}_{g_{i}}=[\Delta x_{ij}\ \ \Delta y_{ij}]^{T}=[x_{g_{i}}-x_{p_{j}}\ \ y_{g_{i}}-y_{p_{j}}]^{T}$, i.e., the vector from the projected coordinate of Gaussian $g_{i}$ $(x_{g_{i}},y_{g_{i}})$ to the coordinate of the current pixel $(x_{p_{j}},y_{p_{j}})$. $\Sigma_{i}=\begin{bmatrix}A_{i}&B_{i}\\ B_{i}&C_{i}\end{bmatrix}$ is the covariance matrix of $g_{i}$.
@@ 3.2. Transformation based on Intra-Tile Coordinates
where $\mathbf{v}{g_{i}}\in\mathbb{R}^{6}$ is a vector determined by Gaussian parameters ($A_{i},B_{i},C_{i}$) and offsets $(\hat{x}{g_{i}},\hat{y}{g_{i}})$, and $\mathbf{v}{p_{j}}\in\mathbb{R}^{6}$ encodes intra-tile relative coordinate terms for pixel $p_{j}$. Notably, the $\vec{\bf{v_{p_{j}}}}$ vectors depend only on intra-tile relative coordinates and constant terms, making them independent of specific Gaussians or tile positions. Hence, $\vec{\bf{v_{p_{j}}}}$ for all pixels in a tile can be precomputed once offline, stored in global memory, and reused throughout the rendering process. In contrast, $\vec{\bf{v_{g_{i}}}}$ must be computed once per Gaussian per tile, as it incorporates both Gaussian parameters and relative offsets $(\hat{x}{g_{i}},\hat{y}{g_{i}})$.
@@ 22        Same as Line 14-21 in Algorithm 1
3.3. GEMM-compatible Blending
@@
(7) ${\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\mathbf{M}_{g}}={\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\begin{bmatrix}\vec{\bf{v}}_{g_{0}}^{T}&\vec{\bf{v}}_{g_{1}}^{T}&\dots&\vec{\bf{v}}_{g_{i}}^{T}&\dots&\vec{\bf{v}}_{g_{255}}^{T}&\end{bmatrix}^{T}}$
${\color[rgb]{0.28515625,0.31640625,0.60546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.28515625,0.31640625,0.60546875}\mathbf{M}_{p}}={\color[rgb]{0.28515625,0.31640625,0.60546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.28515625,0.31640625,0.60546875}\begin{bmatrix}\vec{\bf{v}}_{p_{0}}\ \vec{\bf{v}}_{p_{1}}\ \dots\ \vec{\bf{v}}_{p_{i}}\ \dots\ \vec{\bf{v}}_{p_{255}}\ \end{bmatrix}}$
@@
where $\mathbf{M}_{power}$ is a $256\times 64$ matrix containing all powerijpower{ij} values for Gaussian–pixel pairs within the tile. As discussed in Section 3.2, $\mathbf{M}_{p}$ remains identical across tiles due to fixed intra-tile coordinates, enabling offline precomputation and reuse with negligible overhead.
Algorithm 2 summarizes the blending procedure. In contrast to vanilla 3DGS blending (Algorithm 1), which evaluates $power_{ij}$ per pixel individually, our method directly consumes the offline-precomputed $\mathbf{M}_{p}$. For each tile, the steps are: 1) Fetch Gaussian attributes ($A_{i}$, $B_{i}$, $C_{i}$, $x_{g_{i}}$, $y_{g_{i}}$) for the current Gaussian batch. 2) Compute $\vec{v_{g_{i}}}$ for each Gaussian $\mathcal{G}i$ using dxicdx{ic} and $\hat{y}_{g_{i}}$, forming $\mathbf{M}_{g}$. 3) Perform matrix multiplication (Eq.(8)) to obtain $\mathbf{M}_{power}$. 4) Execute volume rendering, computing $\alpha_{ij}$ using exponentials of $\mathbf{M}_{power}$ values, consistent with Algorithm 1.

### 06 Apr 2026 v2

@@ 2. Background
where $\mathbf{C}_{j}$ is the rendered color of pixel $p_{j}$, and $T_{i}$ represents the accumulated transmittance up to Gaussian ii.
@@ 8    Fetch data of Gaussian $\mathcal{G}_{i}$:
10      covariance matrix: $\Sigma_{i}^{-1}=\begin{bmatrix}A_{i}&B_{i}\\ B_{i}&C_{i}\end{bmatrix}$
@@ 3.1. Calculation Form of $\alpha$’s Exponential
where $\mathbf{x}_{g_{i}}=[\Delta x_{ij}\ \ \Delta y_{ij}]^{T}=[x_{g_{i}}-x_{p_{j}}\ \ y_{g_{i}}-y_{p_{j}}]^{T}$, i.e., the vector from the projected coordinate of Gaussian $g_{i}$ $(x_{g_{i}},y_{g_{i}})$ to the coordinate of the current pixel $(x_{p_{j}},y_{p_{j}})$. $\Sigma_{i}^{-1}=\begin{bmatrix}A_{i}&B_{i}\\ B_{i}&C_{i}\end{bmatrix}$ is the covariance matrix of $g_{i}$.
@@ 3.2. Transformation based on Intra-Tile Coordinates
where ${\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\vec{\bf{v_{g_{i}}}}}\in\mathbb{R}^{6}$ is a vector determined by Gaussian parameters ($A_{i},B_{i},C_{i}$) and offsets $(\hat{x}_{g_{i}},\hat{y}_{g_{i}})$, and ${\color[rgb]{0.28515625,0.31640625,0.60546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.28515625,0.31640625,0.60546875}\vec{\bf{v_{p_{j}}}}}\in\mathbb{R}^{6}$ encodes intra-tile relative coordinate terms for pixel $p_{j}$. Notably, the $\vec{\bf{v_{p_{j}}}}$ vectors depend only on intra-tile relative coordinates and constant terms, making them independent of specific Gaussians or tile positions. Hence, $\vec{\bf{v_{p_{j}}}}$ for all pixels in a tile can be precomputed once offline, stored in global memory, and reused throughout the rendering process. In contrast, $\vec{\bf{v_{g_{i}}}}$ must be computed once per Gaussian per tile, as it incorporates both Gaussian parameters and relative offsets $(\hat{x}_{g_{i}},\hat{y}_{g_{i}})$.
@@ 22        Same as Line 14-21 in Algorithm 1
$3.$ GEMM-compatible Blending
@@
(7) ${\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\mathbf{M}_{g}}={\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\begin{bmatrix}\vec{\bf{v}}_{g_{0}}^{T}&\vec{\bf{v}}_{g_{1}}^{T}&\dots&\vec{\bf{v}}_{g_{i}}^{T}&\dots&\vec{\bf{v}}_{g_{255}}^{T}\end{bmatrix}^{T}}$
${\color[rgb]{0.28515625,0.31640625,0.60546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.28515625,0.31640625,0.60546875}\mathbf{M}_{p}}={\color[rgb]{0.28515625,0.31640625,0.60546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.28515625,0.31640625,0.60546875}\begin{bmatrix}\vec{\bf{v}}_{p_{0}}&\vec{\bf{v}}_{p_{1}}&\dots&\vec{\bf{v}}_{p_{i}}&\dots&\vec{\bf{v}}_{p_{255}}\end{bmatrix}}$
@@
where $\mathbf{M}_{power}$ is a $256\times 256$ matrix containing all $power_{ij}$ values for Gaussian–pixel pairs within the tile. As discussed in Section 3.2, $\mathbf{M}_{p}$ remains identical across tiles due to fixed intra-tile coordinates, enabling offline precomputation and reuse with negligible overhead.
Algorithm 2 summarizes the blending procedure. In contrast to vanilla 3DGS blending (Algorithm 1), which evaluates $power_{ij}$ per pixel individually, our method directly consumes the offline-precomputed $\mathbf{M}_{p}$. For each tile, the steps are: 1) Fetch Gaussian attributes ($A_{i}$, $B_{i}$, $C_{i}$, $x_{g_{i}}$, $y_{g_{i}}$) for the current Gaussian batch. 2) Compute $\color[rgb]{0.6015625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0}\vec{\bf{v_{g_{i}}}}$} for each Gaussian $\mathcal{G}i$ using dxicdx{ic} and $\hat{y}_{g_{i}}$, forming $\mathbf{M}_{g}$. 3) Perform matrix multiplication (Eq.(8)) to obtain $\mathbf{M}_{power}$. 4) Execute volume rendering, computing $\alpha_{ij}$ using exponentials of $\mathbf{M}_{power}$ values, consistent with Algorithm 1.